OpenStudy (anonymous):

(-5) x ( ) + (-5) x ( ) +(2) x ( ) +(2) x ( ) What property is this?

OpenStudy (anonymous):

Distributive \[-5(2+3)+2(2+3)\]would look like \[(-5)\times(2)+(-5)\times(3)+(2)\times(2)+(2)\times(3)\] and without anything in the parenthesis it would look like\[-5(number +number )+2(number + number )\] which would turn into; \[(-5)\times(number)+(-5)\times(number)+(2)\times(number)+(2)\times(number)\]